OpenStudy (anonymous):

Which is the graph of the line with equation: y – 6 = 4(x – 3) + 2

OpenStudy (anonymous):

here are the choices

OpenStudy (solomonzelman):

y – 6 = 4(x – 3) + 2 I would put it into a slope intercept form - y=mx+b. y – 6 = 4(x – 3) + 2 y – 6 = 4x – 12 + 2 y = 4x – 4 plug in points for x, if x=1 then y = 4(1) – 4 y=4-4 y=0 so the first coordinate is going to be (1,0) if x=2 then y = 4(2) – 4 y = 8 – 4 y=4 so the second coordinate is going to be (2,4) graph these 2 points and connect them with a line.